Say true or false and justify your answer:   2^3 × 3^2 = 6^5

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let us consider LHS = 2^3× 3^2
Expansion of 2^3× 3^2 = 2 × 2 × 2 × 3 × 3
= 72
Now, consider RHS = 6^5
Expansion of 6^5= 6 × 6 × 6 × 6 × 6
= 7776
By comparing LHS and RHS,
LHS < RHS
2^3< 5^2
Hence, the given statement is false.

Thanks